OpenStudy (anonymous):

If x is not a whole number times π, then csc(-x) + csc(x) can be simplified as.........

hartnn (hartnn):

use this : \(\csc (-x) = -\csc x\)

OpenStudy (anonymous):

im not sure what to do

hartnn (hartnn):

since, \(\sin (-x) = -\sin x\) and sin x = 1/csc x we also have \(\csc (-x) = - \csc x\) so just replace csc (-x) in your question by - csc x and see what u get ??

OpenStudy (anonymous):

-csc(x) =-cscx

hartnn (hartnn):

csc(-x) + csc(x) = - csc x + csc x = 0 makes sense ?

OpenStudy (anonymous):

oh, yes